OpenStudy (anonymous):

Explain an error pattern in each of the following: a. 13 = 1 , 27 = 2, 16 = 1 55 5 73 3 64 4 b. 4 + 2 = 6, 2 + 3 = 5, 7 + 1 = 8 5 3 8 5 4 9 8 3 11 c. 8 3/8- 6 ¼ = 2 2/4, 5 3/8- 2 2/3 = 3 1/5, 2 2/7 – 1 1/3 = 1 1/4 d. 2/3 . 3 = 6/9, ¼ . 6 = 6/24, 4/5. 2 = 8/10

OpenStudy (turingtest):

a. 13 = 1 , 27 = 2, 16 = 1 55 5 73 3 64 4 Not sure, there seem to be a number of errors here(like dividing the top and bottom by different numbers, which is a no-no), except in the last which is correct. b. 4 + 2 = 6, 2 + 3 = 5, 7 + 1 = 8 5 3 8 5 4 9 8 3 11 The problem here is thinking you can add the denominator straight across. You can't. You must get both fraction over the same denominator first. c. 8 3/8- 6 ¼ = 2 2/4, 5 3/8- 2 2/3 = 3 1/5, 2 2/7 – 1 1/3 = 1 1/4 Again we have the same problem as above, someone is just adding the numbers as they are: \[1/4+1/5\neq1/9\]so that's no good, we need to fix the denominators first. d. 2/3 . 3 = 6/9, ¼ . 6 = 6/24, 4/5. 2 = 8/10 here we seem to have an error in multiplication, someone is multiplying both the numerator and denominator by the number next to it. That's not how this is done. It's done like so\[(1/4)*6=6/4=3/2\]we should only be multiplying the numerator here.